If 8 kg of air at 27°C are heated at constant pressure until the absolute temperature doubles, calculate the heat required in kJ.    


1
Expert's answer
2021-10-18T11:01:51-0400

In absolute (Kelvin) scale, 27oC = 27 + 273 = 300 K.

Therefore, doubling the temperature means increasing it by another 300 K ("\Delta{T}=300K" ).

The specific heat capacity of air at constant pressure (cp) at 300 K is approximately 1.01 "\frac{kJ}{kg\cdot{K}}" .

Now the heat can be calculated:

"Q=c_pm\Delta{T}=1.01\frac{kJ}{kg\cdot{K}}\times8kg\times300K=2424kJ"


Answer: 2424 kJ